OpenStudy (anonymous):

The variables a,b, and c are all integers and a+b+c=50. If a<0 and 0

OpenStudy (anonymous):

I need help asap!>.<

OpenStudy (anonymous):

well, you know that a will be negative...since it is asking for the minimum value of c i would try using -1 for a. 29 for b and then solving for c....that is if i read the question right!

OpenStudy (johnweldon1993):

@Hope_nicole correct :)

OpenStudy (anonymous):

The answer choices are 16, 17,21, 22, and 25

OpenStudy (johnweldon1993):

because you want the minimum that C must be....so use the MAXIMUM that a and b can be...A can be -1...and B can be 29 so if -1 + 29 + c = 50 then 28 + c = 50 c = 22 :)

OpenStudy (anonymous):

well if you plug in the number i said earlier you would have -1+29+c=50 28+c=50 c=22!

OpenStudy (anonymous):

yay! i got it^-^ thanks guys!♥

OpenStudy (anonymous):

The minimum value of c occur when a and b have the greatest value. We know that they are all integers, and b is in the interval 0 < b < 30, so the largest value for b is 29. c is in the interval, a < 0, so the largest value for a is -1. Now that we have the greatest values for b and a, we plug these in and solve for the smallest value of a. a + (29) + (-1) = 50 a = 50 - 29 + 1 a = 22 Therefore, the smallest value for is 22.

OpenStudy (johnweldon1993):

anytime :)

OpenStudy (anonymous):

@Thatyoungbuttercup

OpenStudy (anonymous):

lemme know if you need anymore help :)